# Evaluation of tungsten influx rate using line emissions from W5+ ions in EAST Tokamak

### v1

@@ Abstract
The S/XB ratios (ionization per emitted photon) allow one to relate spectroscopic emissivity measurements to the impurity influx from a localized source. In this work, we determine the tungsten influx by examining two dominant EUV (Extreme Ultraviolet) line emissions at 282.13 Å and 394.07 Å, corresponding to the $4f^{14}5f\rightarrow 4f^{14}5d$ radiative transitions of the W5+ ion. The ground configuration of W5+ consists of the ground level and a metastable level, with the latter having a higher population than the ground state. Therefore, a simple approach assuming that the transitions are independent, i.e., only populated by a unique level source, requires correction.
To address this, we have developed a fully collisional–radiative modeling in which 430 levels contribute to the ionization. We have utilized three advanced computational codes – hullac (Hebrew University - Lawrence Livermore Atomic Code), autostructure, and fac (Flexible Atomic Code) – for the atomic structure calculations. These codes provide the necessary information such as wavelengths, collisional and radiative transition rate coefficients. The fac code was also used to calculate the direct electron-impact ionization under the distorted-wave approximation. We also included contributions to total ionization from excitation-autoionization processes up to $n=15$ manifolds from the distorted-wave calculations.
Subsequently, we used these results to ascertain the tungsten impurity influx in a dedicated discharge of the EAST tokamak, which operates with full tungsten divertors. In our findings, we observed that for the density range relevant to the edge region of a tokamak reactor, the S/XB ratios are almost independent of electron density but exhibit significant variation with electron temperature.
@@ 1 INTRODUCTION
Various intrinsic impurities are inevitably generated in tokamak discharges due to the interaction between plasma and wall. Impurities ionized in the core release a large number of electrons, severely diluting the concentration of the primary plasma and affecting the reaction power density of fusion, decreasing the overall performance of the plasma. Therefore, monitoring and controlling the entry of impurities into the core is an urgent problem for long-pulse discharge operations.
@@
Tungsten has been selected as the divertor and first wall material for ITER due to its high melting point, high thermal conductivity, low sputtering rate, low fuel retention rate, and low neutron activation rate [6]. However, highly ionized tungsten ions in the core and edge emit intense radiation at various wavelength ranges, which can cause significant radiation power loss and severely reduce plasma confinement performance. Therefore, evaluating the amount of tungsten entering through the edge is crucial for functioning high confinement long pulse plasma. Many extensive studies have been dedicated to the evaluation of the tungsten influx using the 4009 Å spectroscopic line of W I [7, 8, 9]. Other lines from W I have been used by Beigman et al. [10], and W II was used by Pospieszczyk et al. [11]. Ballance et al. [12] conducted a detailed study on the S/XB ratios for the 1099.05 Å, 1119.7 Å, 1172.47 Å, and 1186.17 Å lines of W3+, assuming these lines as independent. Dong et al. [13] calculated the influx rate of W6+ using the S/XB coefficients calculated with the ADAS code [14].
@@ 4.1 Atomic Structure
The ion $W^{5+}$ is isoelectronic to the Thulium, having a ground configuration [[[[Xe]]]] $4f^{14}5d$. This configuration has two levels separated by about 1 eV of energy. The lowest level of the ground configuration is $(4f^{14}5d)_{\frac{3}{2}}$ and the second level is $(4f^{14}5d)_{\frac{5}{2}}$. For most of the relevant electronic densities and temperatures, both of these two levels are significantly occupied and in particular, the population of the $J=\frac{5}{2}$ is higher than those of the ground level. We found two radiative transitions well separated and distinguishable in the experimental spectra, corresponding to the $(4f^{14}5f)_{\frac{5}{2}}\rightarrow(4f^{14}5d)_{\frac{3}{2}}$ transition at $\lambda_{1}=382.13$, and $(4f^{14}5f)_{\frac{7}{2}}\rightarrow(4f^{14}5d)_{\frac{5}{2}}$ transition at $\lambda_{1}=394.07$.

### v2

@@ Abstract
The S/XB ratios (ionization events per emitted photon) allow one to relate spectroscopic emissivity measurements to the impurity influx from a localized source. In this work, we determine the tungsten influx by examining two dominant EUV (Extreme Ultraviolet) line emissions at 382.13 Å and 394.07 Å, corresponding to the $4f^{14}5f\rightarrow 4f^{14}5d$ radiative transitions of the W5+ ion. The ground configuration of W5+ consists of the ground level and a metastable level, with the latter having a higher population than the ground state. Therefore, a simple approach assuming that the transitions are independent, i.e., only populated by a unique level source, requires correction.
To address this, we have developed a fully collisional–radiative modeling in which 430 levels contribute to the ionization. We have utilized three advanced computational codes – hullac (Hebrew University - Lawrence Livermore Atomic Code), as (AutoStructure), and fac (Flexible Atomic Code) – for the atomic structure calculations. These codes provide the necessary information such as wavelengths, collisional and radiative transition rate coefficients. The fac code was also used to calculate the direct electron-impact ionization under the distorted-wave approximation. We also included contributions to total ionization from excitation-autoionization processes up to $n=15$ manifolds from the distorted-wave calculations.
Subsequently, we used these results to ascertain the tungsten impurity influx in a dedicated discharge of the EAST tokamak, which operates with full tungsten divertors. In our findings, we observed that for the density range relevant to the edge region of a tokamak plasma, the S/XB ratios are almost independent of electron density but exhibit significant variation with electron temperature.
@@ 1 INTRODUCTION
Various intrinsic impurities are inevitably generated in tokamak discharges due to the interaction between plasma and wall. Impurities ionized in the core release a large number of electrons, severely diluting the concentration of the primary plasma and affecting the reaction power density of fusion, decreasing the overall performance of the plasma. Therefore, monitoring and controlling the entry of impurities into the core are an important issue for long-pulse discharge operations.
@@
Tungsten has been selected as the divertor and first wall material for ITER due to its high melting point, high thermal conductivity, low sputtering rate, low tritium retention rate, and low neutron activation rate [6]. However, highly ionized tungsten ions in the core and edge emit intense radiation at various wavelength ranges, which can cause significant radiation power loss and severely reduce plasma confinement performance. Therefore, evaluating the amount of tungsten entering through the edge is crucial for functioning high confinement long pulse plasma. Many extensive studies have been dedicated to the evaluation of the tungsten influx using the 4009 Å spectroscopic line of W I [7, 8, 9]. Other lines from W I have been used by Beigman et al. [10], and W II was used by Pospieszczyk et al. [11]. Ballance et al. [12] conducted a detailed study on the S/XB ratios for the 1099.05 Å, 1119.7 Å, 1172.47 Å, and 1186.17 Å lines of W3+, assuming these lines as independent. Dong et al. [13] evaluated the influx rate of W6+ using the S/XB coefficients calculated with the ADAS code [14].
@@ 4.1 Atomic Structure
The ion $W^{5+}$ is isoelectronic to the Thulium ($Z=69$), having a ground configuration [[[[Xe]]]] $4f^{14}5d$. This configuration has two levels separated by about 1 eV of energy. The lowest level of the ground configuration is $(4f^{14}5d)_{\frac{3}{2}}$ and the second level is $(4f^{14}5d)_{\frac{5}{2}}$. For most of the relevant electronic densities and temperatures, both of these two levels are significantly occupied and in particular, the population of the $J=\frac{5}{2}$ is higher than those of the ground level. We found two radiative transitions well separated and distinguishable in the experimental spectra, corresponding to the $(4f^{14}5f)_{\frac{5}{2}}\rightarrow(4f^{14}5d)_{\frac{3}{2}}$ transition at $\lambda_{1}=382.13$, and $(4f^{14}5f)_{\frac{7}{2}}\rightarrow(4f^{14}5d)_{\frac{5}{2}}$ transition at $\lambda_{1}=394.07$.